Find the first three x-intercepts of the graph of the given function on the positive x-axis.
f(x) = 2 − 4 cos(x +π/3)
a.
b.2pi
c.

Answer by lwsshak3(11628)

Find the first three x-intercepts of the graph of the given function on the positive x-axis.
f(x) = 2 − 4 cos(x +π/3)
set f(x)=0
2 − 4 cos(x +π/3)=0
4 cos(x+π/3)=2
cos(x+π/3)=2/4=1/2
x+π/3=π/3, 5π/3, 7π/3
x=0, 4π/3, 2π